OpenStudy (anonymous):

the length of a rectangle is three times the width. the area is 108 centimeter square. find the dimensions of the rectangle.

OpenStudy (ikram002p):

 thus area= 3x*x=3x^2=108 3x^2 =108 solve for x hehe

OpenStudy (midhun.madhu1987):